OpenStudy (anonymous):

help PLEASE!!!!! find the vertex, focus, and the directrix of the parabola and sketch its graph. x+y^2=0

OpenStudy (mertsj):

So this one is y^2=-4px Which way does it open?

OpenStudy (anonymous):

down. but wait, how do u know to use that particular equation?

OpenStudy (mertsj):

x^2=4py opens up x^2=-4py opens down y2=4px opens to the right y^2=-4px opens to the left Which way does this one open?

OpenStudy (anonymous):

down and to the left?

OpenStudy (mertsj):

It can't be both. You have 4 choices. up, down, right, or left. Pick one

OpenStudy (anonymous):

opens to the left?

OpenStudy (mertsj):